If one liter equals 0.264 gallons, then 5 gallons equals how many liters?
Answer by ptaylor(2198)


1 Liter =0.264 Gal divide each side by 0.264
(1/0.264)Liter=1 gal multiply each side by 5
(5/0.264)Liter=5 gal
0.264=264/1000
5/(264/1000)=
multiply numerator and denominator by 1000/264
5/(264/1000)=5000/264=625/33
So:
5 gal=(625/33) Liters=18.9393 Liters
Actually you can get the answer faster if you divide 5 by 0.264 (Step 3)
Hope this helps---ptaylor